# Bias Amplification: Language Models as Increasingly Biased Media

### v1

@@ Abstract
As Large Language Models (LLMs) become increasingly integrated into various facets of society, a significant portion of online text consequently become synthetic. This raises concerns about bias amplification, a phenomenon where models trained on synthetic data amplify the pre-existing biases over successive training iterations. Previous literature seldom discusses bias amplification as an independent issue from model collapse. In this work, we address the gap in understanding the bias amplification of LLMs with four main contributions. Firstly, we propose a theoretical framework, defining the necessary and sufficient conditions for its occurrence, and emphasizing that it occurs independently of model collapse. Using statistical simulations with weighted maximum likelihood estimation, we demonstrate the framework and show how bias amplification arises without the sampling and functional form issues that typically drive model collapse. Secondly, we conduct experiments with GPT-2 to empirically demonstrate bias amplification, specifically examining open-ended generational political bias with a benchmark we developed. We observe that GPT-2 exhibits a right-leaning bias in sentence continuation tasks and that the bias progressively increases with iterative fine-tuning on synthetic data generated by previous iterations. Thirdly, we explore three potential mitigation strategies: Overfitting, Preservation, and Accumulation. We find that both Preservation and Accumulation effectively mitigate bias amplification and model collapse. Finally, using novel mechanistic interpretation techniques, we demonstrate that in the GPT-2 experiments, bias amplification and model collapse are driven by distinct sets of neurons, which aligns with our theoretical framework.
Bias Amplification: Language Models as Increasingly Biased Media
Ze Wang1,2, Zekun Wu1,2, Jeremy Zhang3, Navya Jain 2 Xin Guan 1, Adriano Koshiyama 1 1Holistic AI, 2University College London, 3Emory University, Correspondence: ze.wang.19@ucl.ac.uk, zekun.wu@holisticai.com
@@ 1 Introduction
Large language models (LLMs) are trained on vast amounts of text scraped from the internet, which plays a crucial role in improving their capabilities, whether through emergent abilities (Wei et al., 2022) or scaling laws (Kaplan et al., 2020). However, as LLMs become more widely integrated into human society—for example, in content creation and summarization in media, academia, and business (Maslej et al., 2024)—concerns are mounting that a significant portion of online text in the future may be generated, either entirely or partially, by LLMs (Peña-Fernández et al., 2023; Porlezza and Ferri, 2022; Nishal and Diakopoulos, 2024). This highlights a significant and underexplored risk: bias amplification, referring to the degradation of fairness within LLMs over self-consuming training loops (Mehrabi et al., 2022; Taori and Hashimoto, 2022), where models progressively amplify the pre-existing biases. This concern initially arises from the tendency of LLMs to learn from biased datasets. For example, Parrish et al. (2022); Wang et al. (2024); Bender et al. (2021) show that LLMs absorb inherent stereotypes, such as racial and gender discrimination, embedded in human-generated text. Additionally, Haller et al. (2023); Rettenberger et al. (2024a) demonstrated that LLMs can be aligned with specific political ideologies by fine-tuning them on biased datasets. Moreover, Wyllie et al. (2024) shows that classifiers trained on synthetic data increasingly favor a particular class label over successive generations, while the shrinking diversity observed by Alemohammad et al. (2023); Hamilton (2024) suggests a risk that certain demographic groups may become progressively underrepresented in the outputs of LLMs.
@@
The amplification of biases has profound societal implications. It can lead to the perpetuation of stereotypes, reinforcement of social inequalities, and the marginalization of underrepresented groups. In the context of political bias, this can influence public opinion, skew democratic processes, and exacerbate polarization. Understanding and mitigating bias amplification is therefore crucial to ensure that LLMs contribute positively to society and do not inadvertently cause harm. Nevertheless, despite the literature on bias amplification in discriminative models, there is a notable lack of comprehensive frameworks and empirical studies specifically addressing bias amplification for LLMs, as shown in Section 2. In contrast to discriminative models, where bias amplification can be attributed to overfitting the dominant features in the training dataset, its occurrence in LLMs may stem from a more nuanced issue. As we will show in Sections 3 and 5, even when the training data is unbiased, the model can still amplify its pre-existing biases during the training process.
In this paper, we seek to fill this research gap by proposing a theoretical framework that establishes and explains the causes of bias amplification in LLMs. Additionally, we perform both statistical simulation and LLM experiments to demonstrate and exemplify the theorem. In summary, our main contributions are as follows:
1. Theoretical Framework: We establish the necessary and sufficient conditions for bias amplification (i.e. Theorem 1). The theorem aids in understanding the cause of bias amplification and in distinguishing between bias amplification and model collapse. We conduct statistical simulations using weighted maximum likelihood estimation to illustrate the theorem (see Section 3.2).
@@
2. Benchmarking Tool: We trained a highly accurate classifier capable of detecting political leaning in long-text content. With this classifier, we offer a benchmark for evaluating political bias in LLMs through open-generation tasks, filling a gap not covered in current bias studies (see Sections 2 and 4.3).
@@
3. Empirical Demonstration and Mitigation Strategies: We demonstrate bias amplification in terms of political bias in GPT-2 using our benchmarking tool: the model exhibits a right-leaning bias in sentence continuation tasks and becomes increasingly right-leaning over successive generations (see Section 5.1). Additionally, we conducted experiments with three potential mitigation strategies, i.e. Overfitting, Preservation, and Accumulation, comparing their effectiveness, and found that some are surprisingly effective (see Section 5.3).
@@
4. Mechanistic Interpretation: Building on our framework, we propose an innovative mechanistic interpretation pipeline that identifies two distinct sets of neurons responsible for bias amplification and model collapse during iterative fine-tuning experiments with GPT-2. We found minimal overlap between these two sets, supporting our theorem that bias amplification can occur independently of model collapse (see Section 5.4).
The rest of the paper is organized as follows: We first discuss related work in Section 2. In Section 3, we present our theoretical framework for bias amplification and conduct a statistical simulation. Section 4 describes the experimental setup for our LLM experiments, including data preparation and model fine-tuning. In Section 5, we present our empirical findings on bias amplification and the effectiveness of mitigation strategies. Finally, in Section 6, we discuss the implications of our work, concluding with limitations in Section 7.
@@ 2 Background and Related Work
Bias Amplification. Bias amplification in machine learning models has been seriously studied in the context of visual recognition tasks. For example, Zhao et al. (2017) found that Conditional Random Fields can exacerbate social biases present in the training data. This study proposed an in-process mitigation approach, employing Lagrangian Relaxation to enforce constraints that ensure the model’s bias performance remains closely aligned with the biases in the training data. Following this, Mehrabi et al. (2022) proposed the concept of bias amplification in feedback loops, where biased models not only amplify the bias present in their training data but also interact with the world in ways that generate more biased data for future models. Xu et al. (2023); Zhou et al. (2024) examined bias amplification in recommendation models, showing how these models reinforce their understanding of mainstream user preferences from training data, leading to an overrepresentation of such preferences in historical data and neglecting rarely exposed items—similar to the concept of sampling error discussed in (Shumailov et al., 2024). Additionally, Wyllie et al. (2024); Taori and Hashimoto (2022) demonstrated that classifiers trained on synthetic data increasingly favor specific class labels over successive generations. Likewise, Ferbach et al. (2024); Chen et al. (2024) observed bias amplification in generative models such as Stable Diffusion, characterized by the overrepresentation of features from the training dataset. Moreover, we conducted a comprehensive review of the literature on model collapse which we put in Appendix L.
Political Biases. LLMs become a new form of "media" that people rely on for global news (Maslej et al., 2024), attracting significant research attention to political biases. Rettenberger et al. (2024b); Shumailov et al. (2024); Feng et al. (2024) explored the bias through voting simulations within the spectrum of German political parties, consistently finding a left-leaning bias in models like GPT-3 and Llama3-70B. Similarly, for the U.S. political landscape, Rotaru et al. (2024); Motoki et al. (2024) identified a noticeable left-leaning bias in ChatGPT and Gemini when tasked with rating news content, evaluating sources, or responding to political questionnaires.
3 Theoretical Framework
In this section, we formalize the conditions under which bias amplification occurs during recursive training on synthetic data. This framework is designed to intuitively illustrate the primary causes of bias amplification. Next, we demonstrate bias amplification and the theorem using Weighted Maximum Likelihood Estimation (WMLE).
@@ 3.1 The Conditions for Bias Amplification
Bias amplification can arise from two key factors. The first is referred to as bias projection. It arises when the bias projection coefficient (explained below) is negative, indicating that the gradient update during training will fully align with the existing bias in the model parameters, thus amplifying it. Consider a fine-tuning process where the pre-trained model parameters $\theta_{t}$ can be decomposed into unbiased and biased components:
@@
If $c_{t}<0$, the gradient update will reinforce the biased component, leading to bias amplification, i.e. $\Delta|\theta_{\text{biased}}|>0$. This occurs because the gradient descent step moves the parameters further in the direction of the existing bias. Intuitively, the bias in the output is influenced by specific neurons, and increasing the weights of these neurons best aligns with the training objective for the given dataset. This does not require the dataset itself to be biased, but rather that it contains features, aside from bias, that align more closely with the functionalities of biased neurons relative to the unbiased neurons considering the loss function.
The second factor is referred to as constraint deficiency. It occurs when the loss or objective function lacks explicit constraints or regularization to counteract bias projection. As a result, the optimization process does not prevent the reinforcement of bias when bias projection is present. To prevent bias amplification theoretically, the loss function can incorporate de-biasing constraints. Consider a modified fine-tuning objective: $\mathcal{L}_{\text{new}}(\theta)=\mathcal{L}_{\text{original}}(\theta)-\lambda\sum_{i=1}^{n}\max(0,c^{i})^{2}+\gamma\sum_{i=1}^{n}\min(0,c^{i})^{2}$, where $c^{i}$ represents the bias projection coefficient for a specific dimension. $\lambda,\gamma\geq 0$ are the Lagrangian multipliers for the two constraints:
(1) $\sum_{i=1}^{n}\max(0,c^{i})^{2}\geq 0$,   i.e. encouraging positive components.
(2) $\sum_{i=1}^{n}\min(0,c^{i})^{2}\leq 0$,  i.e. eliminating negative components.
The absence of such constraints allows bias amplification when bias projection is present. Based on these findings, we establish the following:
Theorem 1 (Condition for Bias Amplification).
Bias amplification, a self-reinforcing process where a model trained on synthetic data amplifies the pre-existing biases from previous training, leading to an increased bias in the data it generates for future training, occurs if both bias projection and constraint deficiency are present. It occurs only if there is bias projection in at least one parameter dimension and constraint deficiency is present.
Alongside understanding the causes of bias amplification, this also implies that bias amplification can occur even without model collapse. Specifically, statistical approximation and functional expressivity errors [^1] We have not discuss functional approximation error here, as it is a general concept that encompasses any structural setting, from learning to objective function, that may lead to suboptimal training outcomes, making it too general for this discussion.—common causes of model collapse (Shumailov et al., 2024)—are unrelated with bias projection and constraint deficiency. This is further supported by our experiments in Sections 3.2 and 5.
3.2 Weighted Maximum Likelihood Estimation (WMLE)
To demonstrate bias amplification in a controlled setting, we consider a statistical estimation cycle using WMLE, devoid of statistical approximation and functional expressivity errors. The mathematical formulation of WMLE is provided in Appendix A. Specifically, we generate a "biased" pre-training dataset $\mathcal{D}_{\text{pre}}$ consisting of $N=10,000$ samples from a "biased" Beta$(\alpha=3,\beta=2)$ distribution, with a mean of $x=0.6$. Maximum likelihood estimation (MLE) is used to estimate the probability density function (pdf) of this beta distribution, resulting in the ’biased’ pre-trained pdf (referred to as $f_{\text{pre}}$). Our goal in the estimation cycle is to estimate the pdf of the true distribution, Beta$(\alpha=2,\beta=2)$, with a mean of $x=0.5$. We generate $N$ samples from the true distribution, referred to as $D_{\text{real}}$, serving as data for the initial round of training. In the first round, we estimate the true pdf using WMLE with $D_{\text{real}}$ and weights derived from $f_{\text{pre}}$, simulating the impact of the biased pre-trained model during the current training. The newly estimated pdf is referred to as $f_{\text{Generation 0}}$. Then, we generate the synthetic data $D_{\text{Generation 0}}$ of size N using $f_{\text{Generation 0}}$. In the second round, we use WMLE with $D_{\text{Generation 0}}$ and weights derived from $f_{\text{Generation 0}}$ to obtain $f_{\text{Generation 1}}$. We then repeat this process, recursively estimating $f_{\text{Generation 2}}$ through $f_{\text{Generation 10}}$, simulating an iterative training process on synthetic data.
Results:
As depicted in Figure 1, the estimated distributions progressively shift towards the mean of the biased Beta$(\alpha=3,\beta=2)$ distribution, at $x=0.6$, becoming increasingly peaked with each generation. This behavior exemplifies bias amplification. The distortion arises from two sources: bias projection, where the estimation process emphasizes data regions where the ’biased’ pre-trained distribution assigns higher probability; and constraint deficiency, where WMLE lacks mechanisms to counteract the bias introduced by the weights, allowing the bias to be amplified over generations. For comparison, Figure 2 shows the results using standard MLE without weighting. The estimated distributions remain consistent across generations, accurately capturing the true Beta(2,2)22(2,2)( 2 , 2 ) distribution. This indicates that without the influence of biased pre-trained pdf and the weighted learning process, bias amplification does not occur.
Figure 1: Weighted Maximum Likelihood Estimation (WMLE) over 10 generations.
Figure 2: Standard Maximum Likelihood Estimation (MLE) over 10 generations, starting from the true distribution Beta(2,2)22(2,2)( 2 , 2 ) without weighting.
4 Experiment Design
@@
In this section, we present and explain in detail the experiments designed to investigate the bias amplification of LLMs (see results in Section 5). Specifically, we analyze the sequential and synthetic fine-tuning of GPT-2. While similar experiments could be conducted with larger models such as GPT-4 or Claude in non-fine-tuning contexts, doing so would likely incur significant environmental costs. The focus of this study is on political bias within the US political spectrum, particularly in the context of sentence continuation tasks. This is important because LLMs are increasingly becoming a source of information for global news consumption Maslej et al. (2024); Peña-Fernández et al. (2023); Porlezza and Ferri (2022), and traditional news outlets, such as the Associated Press, are beginning to incorporate LLMs to automate content generation from structured data (The Associated Press, 2024).
4.1 Experiment Dataset Preparation

### v2

# Bias Amplification: Large Language Models as Increasingly Biased Media

@@ Abstract
Model collapse—a phenomenon where models degrade in performance due to indiscriminate use of synthetic data—is well studied. However, its role in bias amplification—the progressive reinforcement of pre-existing social biases in Large Language Models (LLMs)—remains underexplored. In this paper, we formally define the conditions for bias amplification and demonstrate through statistical simulations that bias can intensify even in the absence of sampling errors, the primary driver of model collapse. Empirically, we investigate political bias amplification in GPT-2 using a custom-built benchmark for sentence continuation tasks. Our findings reveal a progressively increasing right-leaning bias. Furthermore, we evaluate three mitigation strategies—Overfitting, Preservation, and Accumulation—and show that bias amplification persists even when model collapse is mitigated. Finally, a mechanistic interpretation identifies distinct sets of neurons responsible for model collapse and bias amplification, suggesting they arise from different underlying mechanisms.
Bias Amplification: Large Language Models as Increasingly Biased Media
Ze Wang1,2thanks: Corresponding Authors, Zekun Wu1,211footnotemark: 1, Jeremy Zhang3, Xin Guan 1, Navya Jain 2 Skylar Lu 3, Saloni Gupta 4, Adriano Koshiyama 1 1Holistic AI, 2University College London 3Emory University, 4University of Maryland, College Park
@@ 1 Introduction
Large Language models (LLMs) are trained on vast amounts of text scraped from the internet, which plays a crucial role in improving their capabilities, whether through emergent abilities (Wei et al., 2022) or scaling laws (Kaplan et al., 2020). However, as they become more widely integrated into human society—for example, in content creation and summarization in media, academia, and business (Maslej et al., 2024)—concerns are mounting that a significant portion of online text in the future may be generated, either entirely or partially, by LLMs (Peña-Fernández et al., 2023; Porlezza and Ferri, 2022; Nishal and Diakopoulos, 2024). This highlights a significant and underexplored risk: bias amplification, where pre-existing biases become progressively reinforced and intensified as models are repeatedly trained on synthetic data (Mehrabi et al., 2022; Taori and Hashimoto, 2022). This concern initially arises from the tendency of LLMs to learn from biased datasets. For example, Parrish et al. (2022); Wang et al. (2024); Bender et al. (2021) show that LLMs absorb inherent biases embedded in human-generated text. Haller et al. (2023); Rettenberger et al. (2024a) demonstrated that LLMs can be aligned with specific political ideologies by fine-tuning them on biased datasets. Moreover, Wyllie et al. (2024) shows that classifiers trained on synthetic data increasingly favor a particular class label over successive generations, while the shrinking diversity observed by Alemohammad et al. (2023); Hamilton (2024) suggests a risk that certain demographic groups may become underrepresented in the outputs of LLMs.
@@
The amplification of biases has profound societal implications. It can lead to the perpetuation of stereotypes, reinforcement of social inequalities, and the marginalization of underrepresented groups. In the context of political bias, this can influence public opinion, skew democratic processes, and exacerbate polarization. Therefore, understanding and mitigating bias amplification is both critical and urgent. Nevertheless, despite the literature on discriminative models, there is a notable lack of comprehensive frameworks and empirical studies specifically addressing bias amplification in Language Model.
Figure 1: Experimental Procedure.
In this paper, we aim to address this research gap by introducing a framework that explains the underlying causes of bias amplification in LLMs. We validate the framework through both statistical simulations and direct experiments, demonstrating the emergence of bias amplification. Furthermore, our findings show the potential distinction between bias amplification and model collapse. In summary, our key contributions are as follows:
1. Formal Conditions: We formally delineate the conditions for bias amplification (Section 3.1), providing the theoretical basis for understanding its causes and relationship to model collapse. We then validate this framework through statistical simulations leveraging a weighted maximum likelihood estimation cycle (Section 3.2).
@@
2. Political Bias Benchmark: We trained a highly accurate classifier capable of detecting political leaning in long-text content, specifically within the context of the US political spectrum. With this classifier as a metric model, we offer a benchmark for evaluating political bias in LLMs through sentence continuation tasks.
@@
3. Empirical Validation: We evaluate the amplification of political bias in GPT-2, where the model exhibits a right-leaning bias in sentence continuation tasks and shifts further to the right over generations (Section 5.1). Additionally, we experimented with three potential mitigation strategies—Overfitting, Preservation, and Accumulation—comparing their effectiveness in both Model collapse and Bias Amplification (Section 5.3). This experiment setup could be easily extended to larger models or different types of bias.
@@
4. Mechanistic Interpretation: We conduct a mechanistic analysis identifying two distinct sets of neurons responsible for bias amplification and model collapse during iterative synthetic fine-tuning with GPT-2. Our findings reveal only partial overlap between these sets, suggesting that bias amplification and model collapse arise from distinct mechanisms (Section 5.4).
@@ 2 Background and Related Work
Bias Amplification has been studied in various domains Zhao et al. (2017) found that Conditional Random Fields can exacerbate social biases present in the training data. It proposed an in-process mitigation approach, employing Lagrangian Relaxation to enforce constraints that ensure the model’s bias performance remains closely aligned with the biases in the training data. Following this, Mehrabi et al. (2022) proposed the concept of bias amplification in feedback loops, where biased models not only amplify the bias present in their training data but also interact with the world in ways that generate more biased data for future models. Xu et al. (2023); Zhou et al. (2024) examined bias amplification in recommendation models, showing how these models reinforce their understanding of mainstream user preferences from training data, leading to an overrepresentation of such preferences in historical data and neglecting rarely exposed items—similar to the concept of sampling error discussed in (Shumailov et al., 2024). Wyllie et al. (2024); Taori and Hashimoto (2022) demonstrated that classifiers trained on synthetic data increasingly favor specific class labels over successive generations. Likewise, Ferbach et al. (2024); Chen et al. (2024) observed bias amplification in generative models such as Stable Diffusion, characterized by the overrepresentation of features from the training dataset. Also, we provide a comprehensive literature review of model collapse in Appendix J.
(a) Weighted Maximum Likelihood Estimation.
(b) Maximum Likelihood Estimation.
Figure 2: Comparison of WMLE and MLE over 10 generations.
Political Biases. In parallel, growing attention has been paid to political biases in LLMs, now a prevalent form of “media” that people rely on for global news (Maslej et al., 2024). Rettenberger et al. (2024b); Shumailov et al. (2024); Feng et al. (2024) explored the bias through voting simulations within the spectrum of German political parties, consistently finding a left-leaning bias in models like GPT-3 and Llama3-70B. Similarly, for the U.S. political landscape, Rotaru et al. (2024); Motoki et al. (2024) identified a noticeable left-leaning bias in ChatGPT and Gemini when tasked with rating news content, evaluating sources, or responding to political questionnaires.
3 Formal Conditions
In this section, we formalize the conditions for bias amplification, offering an intuitive look at its principal drivers. We then illustrate these ideas using Weighted Maximum Likelihood Estimation (WMLE).
@@ 3.1 The Conditions for Bias Amplification
The primary factor is referred to as bias projection. It arises when the bias projection coefficient is negative. To illustrate this, consider a fine-tuning process in which the pre-trained model parameters $\theta_{t}$ can be expressed as the sum of unbiased and biased components:
@@
If $c_{t}<0$, the gradient update will reinforce the biased component, leading to bias amplification, i.e. $\Delta|\theta_{\text{biased}}|>0$. This occurs because the gradient descent step moves the parameters further in the direction of the existing bias. Intuitively, bias in the output arises from specific neurons, whose weights increase when their activations align with the fine-tuning dataset. If biased neurons capture more relevant patterns than unbiased ones, optimization will further reinforce their weights.
The second is sampling error, akin to statistical approximation error Shumailov et al. (2024). If the model has a pre-existing bias, it inherently assigns higher probabilities to tokens that produce biased outputs. Consequently, during synthetic data generation, unbiased tokens—and thus unbiased samples—are more likely to be lost at each resampling step with a finite sample, though this error vanishes as the sample size approaches infinity. This overrepresents biased patterns in the synthetic data, surpassing the model’s original bias and true next-token probabilities. Sampling error thus complements bias projection by further activating biased neurons in response to the skewed dataset.
By definition, bias projection is a sufficient condition for bias amplification, while sampling error serves as a complementary factor. However, sampling error is a sufficient condition for model collapse to occur with nonzero probability Shumailov et al. (2024). This distinction motivates our investigation into whether bias amplification can occur without model collapse, as this would necessitate additional mitigation strategies beyond those for model collapse.
3.2 Statistical Simulation
To simulate a controlled setting without sampling error, we consider a statistical estimation cycle using WMLE [^1] The mathematical formulation is detailed in Appendix A. with a large sample size of each resampling step. Specifically, we generate a pre-training dataset $\mathcal{D}_{\text{pre}}$ with 100,000 samples from a Beta$(3,2)$ distribution, representing a biased pre-training dataset. Using maximum likelihood estimation (MLE), we estimate its probability density function, yielding the pre-trained model $f_{\text{pre}}$.
Next, we fine-tune $f_{\text{pre}}$ to approximate a different distribution, Beta(2,2)22(2,2)( 2 , 2 ). We generate 100,000 samples from this distribution, denoted as $D_{\text{real}}$, which serves as the initial fine-tuning dataset. In the first round, we apply weighted maximum likelihood estimation (WMLE) using weights derived from $f_{\text{pre}}$, which encode the pre-existing bias of the pre-trained model. This weighting captures the influence of the pre-trained model’s parameters on subsequent training. This produces the fine-tuned model $f_{\text{0}}$. We then generate a synthetic dataset $D_{\text{0}}$ of the same size using $f_{\text{0}}$, initiating the iterative fine-tuning loop. In each subsequent round, WMLE is applied using $D_{\text{k}}$ with weights from $f_{\text{k}}$, resulting in $f_{\text{k+1}}$. This process is repeated iteratively, producing models $f_{\text{1}}$ through $f_{\text{10}}$.
Figure 2(a) shows the estimated distributions gradually shift toward the mean of the biased pre-training dataset at $x=0.6$, becoming progressively more peaked over generations. This occurs despite further training on samples drawn from Beta(2,2)22(2,2)( 2 , 2 ) and synthetic data generated from successive models. The distortion arises because the fine-tuning process disproportionately emphasizes regions where the pre-trained distribution assigns higher probability, leading to biased learning. For comparison, Figure 2(b) presents the results using standard MLE without weighting. In this case, the estimated distributions remain stable across generations, accurately representing the Beta(2,2)22(2,2)( 2 , 2 ) distribution.
4 Experimental Design
@@
This section provides the details of the experiments on LLMs, focusing on the sequential and synthetic fine-tuning of GPT-2 under different setups [^2] Similar experiments with larger models like GPT-4 or Claude can be easily adapted to the current setup but would incur significant environmental costs.. The step-by-step experimental procedure is outlined in Figure 1. Our study focuses on the political bias of LLMs within the US political spectrum, particularly in sentence continuation tasks. This is important as LLMs are increasingly influencing global news consumption Maslej et al. (2024); Peña-Fernández et al. (2023); Porlezza and Ferri (2022), and traditional news outlets, such as the Associated Press, are beginning to integrate LLMs for automated content generation from structured data (The Associated Press, 2024).
4.1 Dataset Preparation